Prove that for Re(s) >1,we have phi(s) =s integral from 1 to infinity f(x)/x^{s+1} dx
1
Expert's answer
2020-09-01T16:44:18-0400

Let "f" and "\phi(s)" be given as in the theorem, and fix "s" with finite abscissa of convergence "\sigma>min(0,\sigma_c)" , so the Dirichlet series

"\phi(s)=\displaystyle\sum_{n=1}^\infty f(n)n^{-s}"

converges at "s". Write

"f(x)=\displaystyle\sum_{n=1}^\infty f(n)\chi(x,n)"

where

"\chi(x,n)={1, if n\geqslant x\brace 0, if n<x}"

Then, for every "X\geqslant 1"

"s\intop_1^X \frac{f(x)}{x^{s+1}}dx=s\intop_1^X\displaystyle\sum_{n=1}^\infty\chi(x,n)\frac{f(n)}{x^{s+1}}dx=" "s\intop_1^X\displaystyle\sum_{n\leqslant X}\chi(x,n)\frac{f(n)}{x^{s+1}}dx="

"=s\displaystyle\sum_{n\leqslant 1}f(n)\intop_1^X\chi(x,n)\frac{1}{x^{s+1}}dx=" "s\displaystyle\sum_{n\leqslant 1}f(n)\intop_1^n\frac{1}{x^{s+1}}dx=" "=\displaystyle\sum_{n\leqslant 1}f(n)\frac{1}{s}(\frac{1}{n^s}-\frac{1}{X^s})=" "\displaystyle\sum_{n\leqslant 1}\frac{f(n)}{n^s}-\frac{f(x)}{X^s}"

Now let "X->\infty" . Then, by the convergence of "\phi(s)", the first term on the right tends to "\phi(s)". Moreover, Kronecker’s Lemma implies that the second term tends to 0. Hence we conclude

"lim_{X->\infty}s\intop_1^X \frac{f(x)}{x^{s+1}}=\phi(s)".

which proves our formula. "\blacksquare"